# FLP-XR: Future Location Prediction on Extreme Scale Maritime Data in Real-time

### v1

@@ IV-D Performance in Resource-Constrained Environments
Finally, to optimize the performance of our architecture, as we mentioned earlier we implemented the entire preprocessing pipeline in Rust, selected for its strong performance across a wide range of computing environments. In order to demonstrate it, we perform the following experiment: for each combination of dataset and hardware environment and having already processed the half of the dataset (for fairness purposes, in order to avoid cold start issues), we feed FLP-XR (see Figure 2) with a batch of records (in two alternative sizes, either 10K or 100K records per batch) and we calculate the time required per record for the preprocessing and the inference step. In Table V, apart from the respective times required for each of the two steps per record (in microseconds), we calculate the end-to-end throughput of our pipeline for the entire batch (in seconds), as the sum of the preprocessing and inference times multiplied by the number of records in the batch. It is clear from these figures that the proposed architecture is able to support a fleet of order 100K (see Xeon+A100 and NUC-i7 rows) or 10K (see RPi 4 rows) vessels in usually less than one or at most few seconds per batch. To put this outcome into a real use-case perspective, in a cloud environment, an AIS monitoring system with world-wide coverage tracks approximately 270K vessels concurrently [^6] Indicatively, see the statistics of the MarineTraffic platform (https://www.marinetraffic.com)., which transmit their position every few seconds while underway [^7] The current AIS protocol guidelines define a 2 to 10 seconds transmission rate of Class A AIS positions while a vessel is underway (and every 3 minutes while at anchor); source: https://www.navcen.uscg.gov/ais-class-a-reports., whereas in an edge environment, an AIS antenna located in a busy port tracks around 500 vessels [^8] Indicatively, see the statistics of the AIS stations at MarineTraffic (URL: https://www.marinetraffic.com/en/ais/details/stations) in Rotterdam (ID:2538) and in Singapore (ID:17406).. As it turns out from the findings of Table V, with the appropriate selection of hardware settings, our methodology is expected to be capable of providing real-time support under all these specs in the Computing Continuum.

### v2

@@ IV-D Performance in Resource-Constrained Environments
Finally, to optimize the performance of our architecture, as we mentioned earlier we implemented the entire preprocessing pipeline in Rust, selected for its strong performance across a wide range of computing environments. In order to demonstrate it, we perform the following experiment: for each combination of dataset and hardware environment and having already processed the half of the dataset (for fairness purposes, in order to avoid cold start issues), we feed FLP-XR (see Figure 2) with a batch of records (in two alternative sizes, either 10K or 100K records per batch) and we calculate the time required per record for the preprocessing and the inference step. In Table V, apart from the respective times required for each of the two steps per record (in microseconds), we calculate the end-to-end throughput of our pipeline for the entire batch (in seconds), as the sum of the preprocessing and inference times multiplied by the number of records in the batch. It is clear from these figures that the proposed architecture is able to support a fleet of order 100K (see Xeon+A100 and NUC-i7 rows) or 10K (see RPi 4 rows) vessels in usually less than one or at most few seconds per batch. To put this outcome into a real use-case perspective, in a cloud environment, an AIS monitoring system with world-wide coverage tracks approximately 270K vessels concurrently [^6] Indicatively, see the statistics of the MarineTraffic platform (https://www.marinetraffic.com)., which transmit their position every few seconds while underway [^7] The current AIS protocol guidelines define a 2 to 10 seconds transmission rate of Class A AIS positions while a vessel is underway (and every 3 minutes while at anchor); source: https://www.navcen.uscg.gov/ais-class-a-reports., whereas in an edge environment, an AIS antenna located in a busy port could potentially track around 2K vessels [^8] Indicatively, see the statistics about the vessel density in Rotterdam area (https://www.marinetraffic.com/en/details/areas/areaId:12/) and Singapore area (https://www.marinetraffic.com/en/details/areas/areaId:112).. As it turns out from the findings of Table V, with the appropriate selection of hardware settings, our methodology is expected to be capable of providing real-time support under all these specs in the Computing Continuum.